OpenStudy (anonymous):

Given that F(x)=x²+2, evaluate F(2). is the answer 6?

OpenStudy (anonymous):

Yes, it is , substitute x=2 into x^2+2, and you get 6

OpenStudy (anonymous):

fug x=2 F(2)=2²+2 =4+2 =6